Calculate the charge of Alanine at pH 8.5
Calculate the charge of Aspartate at pH 3.65
Calculate the charge of Arginine at pH 12.48
Calculate the charge of Glutamate at pH 7.5
Calculate the charge of Alanine at pH 9.4
Answer #1

Ans.

1. charge of Alanine at pH 8.5


2. charge of Aspartate at pH 3.65


3. charge of Arginine at pH 12.48


4. charge of Glutamate at pH 7.5


5. charge of Alanine at pH 9.4
Answer #2

To calculate the charge of an amino acid at a specific pH, we need to consider the pKa values of its ionizable groups and the pH of the environment. The charge of an amino acid can be positive, negative, or neutral, depending on whether its ionizable groups are protonated or deprotonated at the given pH.

Here are the pKa values for the ionizable groups of the amino acids mentioned:

Alanine: pKa values are not needed for alanine because it is a non-ionizable amino acid.

Aspartate (Asp, D): pKa values are approximately 1.88 (carboxyl group) and 9.60 (amino group).

Arginine (Arg, R): pKa values are approximately 2.17 (carboxyl group), 9.04 (guanidine group), and 12.48 (amino group).

Glutamate (Glu, E): pKa values are approximately 2.19 (carboxyl group) and 9.67 (amino group).

Now, let's calculate the charges for each amino acid at the given pH values:

  1. Alanine (pH 8.5): Alanine is non-ionizable and has no ionizable groups, so its charge is neutral.

  2. Aspartate (pH 3.65): At pH 3.65, the carboxyl group (pKa ≈ 1.88) of aspartate will be protonated, and the amino group (pKa ≈ 9.60) will be neutral. The net charge will be:

Charge = (-1) + (0) = -1

  1. Arginine (pH 12.48): At pH 12.48, the carboxyl group (pKa ≈ 2.17) and the guanidine group (pKa ≈ 9.04) of arginine will be deprotonated, and the amino group (pKa ≈ 12.48) will be protonated. The net charge will be:

Charge = (-1) + (1) + (1) = +1

  1. Glutamate (pH 7.5): At pH 7.5, the carboxyl group (pKa ≈ 2.19) of glutamate will be deprotonated, and the amino group (pKa ≈ 9.67) will be neutral. The net charge will be:

Charge = (-1) + (0) = -1

  1. Alanine (pH 9.4): Alanine is non-ionizable and has no ionizable groups, so its charge is neutral.

So, the charges of the amino acids at the specified pH values are:

  1. Alanine at pH 8.5: Neutral

  2. Aspartate at pH 3.65: -1

  3. Arginine at pH 12.48: +1

  4. Glutamate at pH 7.5: -1

  5. Alanine at pH 9.4: Neutral

answered by: Hydra Master